# PiCO: Peer Review in LLMs based on the Consistency Optimization

### v2

@@ Abstract
Existing large language models (LLMs) evaluation methods typically focus on testing the performance on some closed-environment and domain-specific benchmarks with human annotations. In this paper, we explore a novel unsupervised evaluation direction, utilizing peer-review mechanisms to measure LLMs automatically without any human feedback. In this setting, both open-source and closed-source LLMs lie in the same environment, capable of answering unlabeled questions and evaluating each other, where each LLM’s response score is jointly determined by other anonymous ones. To obtain the ability hierarchy among these models, we assign each LLM a learnable capability parameter to adjust the final ranking. We formalize it as a constrained optimization problem, intending to maximize the consistency of each LLM’s capabilities and scores. The key assumption behind is that high-level LLM can evaluate others’ answers more accurately than low-level ones, while higher-level LLM can also achieve higher response scores. Moreover, we propose three metrics called PEN, CIN, and LIS to evaluate the gap in aligning human rankings. We perform experiments on multiple datasets with these metrics, validating the effectiveness of the proposed approach.
@@ 1 Introduction
Large language models (LLMs)[11; 2; 12; 47] have achieved remarkable success across a variety of real-world applications [58; 36; 40; 56]. With the increasingly widespread application of these models, there is an urgent need for an effective evaluation method to ensure that their performance and usability meet the growing demands. To assess the ability level of LLMs, a large number of evaluation benchmarks have been proposed by using some small and domain-specific datasets with human-curated labels, such as MMLU [28], HELM [34], Big-Bench[43], GLUE[49]. However, these benchmarks can only measure LLMs’ core capability on a confined set of tasks (e.g. multi-choice knowledge or retrieval questions), which fails to assess their alignment with human preference in open-ended tasks adequately [16; 31; 38]. On the other hand, these evaluations may suffer from benchmark leakage issue, referring that the evaluation data is unknowingly used for model training, which can also lead to misleading evaluations [53; 60]. Therefore, blindly improving scores on these public benchmarks cannot always yield a large language model that truly satisfies human requirements.
@@
For assessing human preferences, recent studies have focused on building crowdsourced battle platforms with human ratings as the primary evaluation metric. Typical platforms include Chatbot Arena [59], MT-Bench [59], and Alpaca-Farm [22]. It constructs anonymous battles between chatbots in real-world scenarios, where users engage in conversations with two chatbots at the same time and rate their responses based on personal preferences. While human evaluation is the gold standard for measuring human preferences, it is exceptionally slow and costly[59]. In addition, adding a new LLM to the crowdsourced battle platforms also poses a cold-start issue [15]. Thus, a fundamental question arises: can we construct an unsupervised LLMs evaluation system without relying on any human feedback?
@@
Actually, in real human evaluation systems, people build their ability hierarchy based on different empirical assumptions. For example, majority voting [23; 10; 44] and rating voting [5] methods are widely used during the decision-making process, which are based on the wisdom of the crowds [44; 13; 54] and have been proven to lead to better results than that of an individual. Moreover, in the established practice of peer-review in academic research, scholars evaluate their academic level rankings based on the consistency assumption, i.e., scholars with stronger abilities have stronger persuasiveness for evaluating others, and can also obtain higher achievements. This paper attempts to explore whether similar phenomena exist in the LLMs evaluation systems.
@@
Figure 1: The framework of PiCO. In this framework, both open-source and closed-source LLMs lie in the same environment, capable of answering unlabeled questions and evaluating each other, where each LLM’s response score is jointly determined by other anonymous ones. We assign each LLM a learnable capability weight to optimize the score ranking based on the consistency assumption, while reducing the entropy of the peer-review evaluation system. The consistency optimization aims to find a final score ranking that all LLMs “agree” it.
@@
In this work, we propose PiCO, a Peer review approach in LLMs based on Consistency Optimization. In this setting, LLMs themselves act as “reviewers”, engaging in mutual assessments to achieve comprehensive, efficient, and performance evaluations without relying on manually annotated data. This method aims to address the limitations of existing evaluation approaches and provide insights into LLMs’ real-world capabilities. As shown in Figure 1, both open-source and closed-source LLMs lie in the same environment and answer the open-ended questions from an unlabeled dataset. Then, we construct anonymous answer pairs, while randomly selecting other LLMs as “reviewers” to evaluate both responses with a learnable confidence weight $w$. Finally, we employ this weight and calculate the response scores $G$ for each LLM based on the weighted joint evaluation. It is worth noting that the whole peer-review process works in an unsupervised way, and our goal is to optimize the confidence weights that re-rank the LLMs to be closer to human rankings.
@@
To evaluate the gap in aligning human rankings, we propose three metrics called PEN (Permutation Entropy), CIN (Count Inversions), LIS (Longest Increasing Subsequence). The experiments are conducted on multiple crowdsourcing datasets and validated on these three metrics. The experimental results demonstrate that the proposed PiCO framework can effectively obtain a large language models’ leaderboard closer to human preferences.
The contributions of this paper can be summarized as follows.
  • We explore a novel unsupervised LLM evaluation direction without human feedback, utilizing peer-review mechanisms to measure LLMs automatically. All LLMs can answer unlabeled questions and evaluate each other.
@@
  • We propose three metrics called PEN, CIN, and LIS on the PiCO framework for evaluating the gap with human preferences.
  • The experiments with these metrics on three crowdsourcing datasets validate the effectiveness of the proposed approach.
@@ 2 The Proposed Approach
In this section, we first describe the problem definition and preference alignment evaluation, and then introduce the proposed PiCO framework in detail.
2.1 Definition and Metrics
Problem Definition. In this subsection, we aim to measure the ability level of LLMs automatically without relying on human annotations. Thus we consider an unsupervised LLM evaluation scenario with an unlabeled dataset $\mathcal{Q}$ consisting of $n$ open-ended questions, where $\mathcal{Q}=\{Q_{i}\}_{i=1}^{n}$. In addition, we have a large language model candidate pool =$\mathcal{M}=\{M_{j}\}_{j=1}^{m}$, which includes both open-source and closed-source models. Write ${M}_{1}\succ{M}_{2}$ to indicate that the LLM ${M}_{1}$ has stronger capabilities than the LLM ${M}_{2}$. Thus, we can assume that the ground-truth ranking $\mathcal{R}^{*}$ alignment with human preferences as follows,
$\mathcal{R}^{*}:=[{M}_{1}\succ{M}_{2}\succ{M}_{3}\succ...\succ{M}_{m}],$ (1)
and assume that the learned ranking $\hat{\mathcal{R}}$ by different evaluation methods is as follows,
@@ 2.1 Definition and Metrics
The goal is to build an LLM ranking $\hat{\mathcal{R}}$ that aligns with human ranking $\mathcal{R}^{*}$, making the loss $\mathcal{L}$ of the both rankings tend towards 00, i.e., (^$\mathcal{L}(\hat{\mathcal{R}},\mathcal{R}^{*})\rightarrow 0$
Preference Alignment Metrics. Before building LLM rankings, we first need to discuss how to evaluate aligned human rankings. Intuitively, the metrics we want mainly describe the differences between two arrays composed of ranking indices. Assuming that human ranking $\mathcal{R}^{*}$ is defined as being well-ranked in ascending order ($[1,2,3,...,m]$) as shown in Eq 1. Thus the metric is to quantify the randomness of the learned ranking array ($[3,1,2,...,m]$) as shown in Eq 2. Based on this, we propose three metrics called PEN, CIN, and LIS, respectively.
PEN (Permutation Entropy). Permutation entropy [8] is a concept used to quantify the complexity or randomness of time series data. It provides a measure of the irregularity or unpredictability of the order of values in a sequence. We thus utilize it to measure the gap with human rankings as follows,
$\mathcal{L}_{PEN}(\hat{\mathcal{R}},\mathcal{R}^{*}):=-\sum p(\pi)\log p(\pi),$ (3)
where
Figure 2: Preference alignment metric. Three metrics for evaluating the gap with human preferences called PEN, CIN, and LIS, respectively
CIN (Count Inversions). Counting inversions [30] aims to measure the degree of disorder or "invertedness" in an array or sequence of elements. We thus define it as follows,
$\mathcal{L}_{CIN}(\hat{\mathcal{R}},\mathcal{R}^{*}):=\sum_{M_{i},M_{j}\sim\mathcal{M}}\mathbf{1}\{M_{i}\succ M_{j}\land i<j\}.$ (4)
LIS (Longest Increasing Subsequence). The longest increasing subsequence aims to find the length of the longest subsequence in a given sequence of elements, where the subsequence is in increasing order. We utilize it to measure the degree of match with human rankings as follows,
$\mathcal{L}_{LIS}(\hat{\mathcal{R}},\mathcal{R}^{*}):=\max\left\{dp[i]\mid 1\leq i\leq m\right\}$ (5)
where
$dp[i]=1+\max\left\{dp[j]\mid 1\leq j<i\land M_{j}\prec M_{i}\right\}.$
@@ 2.2 Algorithm Details
The PiCO framework, depicted in Figure 3, involves peer-review and consistency optimization stages. In the peer-review stage, we first collect an unlabeled dataset $\mathcal{Q}$ consisting of open-ended questions, and construct a large language model pool $\mathcal{M}$ that includes both open-source and closed-source LLMs. Then, we let all LLMs answer each unlabeled question to obtain the response set $\mathcal{A}$. We shuffle the set and construct anonymous answer pairs, while randomly selecting other LLMs as “reviewers” to evaluate both responses with a learnable confidence $w$. Finally, we can obtain the answer-ranking data $\mathcal{D}$ and calculate the response score $G$ for each large language model. In the consistency optimization phase, we maximize the consistency of each LLM’s capability $w$ and score $G$ with constrained optimization, while re-ranking the LLMs to be closer to human rankings.
Figure 3: The pipeline of the PiCO. It is mainly composed of two components: the peer-review and consistency optimization stages. Specifically, in the peer-review stage, the unlabeled dataset $\mathcal{Q}$ and the LLMs candidate pool $\mathcal{M}$ are given. Then, we let all LLMs answer each unlabeled question to obtain the response set $\mathcal{A}$. We shuffle the set and construct anonymous answer pairs, while randomly selecting other LLMs to evaluate both responses with a learnable confidence $w$. As a result, we can obtain the answer-ranking data $\mathcal{D}$ which is a quadruple that records the partial order between two answers and the evaluator’s confidence weight. In the consistency optimization stage, we update the parameter $w$ by maximizing the consistency of each LLM’s capability and score, while re-ranking the LLMs to be closer to human rankings.
2.2.1 Peer Review Stage
Data Collection and LLMs Candidate Pool Construction. Benefiting from the creation of crowdsourced battle platforms, we accessed open assessment datasets from Chatbot Arena[59], MT-Bench[59], and AlpacaEval[32]. These open datasets include critical fields such as "question__\__id" and "question__\__content." Utilizing the Chatbot Arena dataset, which features pairwise data from twenty LLMs with human preference annotations, we assembled an LLM pool =$\mathcal{M}=\{M_{j}\}_{j=1}^{m}$. Leveraging 33K human-annotated interactions from this dataset, we established a ground-truth ranking $\mathcal{R}^{*}$ and gathered responses $\mathcal{A}=\{\{A_{i}^{j}\}_{i=1}^{n}\}_{j=1}^{m}$ for our dataset $\mathcal{Q}=\{Q_{i}\}_{i=1}^{n}$.
Answer-Ranking Data Construction Based on Peer Review. After obtaining the responses set $\mathcal{A}$, we aim to generate answer-ranking data $\mathcal{D}$ through the peer-review mechanism. Specifically, for the same question $Q_{i}\in\mathcal{Q}$, we randomly construct a battle pair $<A_{i}^{j},A_{i}^{k}>$ for review. Each battle pair will be randomly assigned five models (“reviewers”) to determine the winners or declare ties. Note that the model may evaluate its own answers, but the entire process is anonymous. As a result, we can obtain the quadruples $(A^{j}_{i},A^{k}_{i},>w^{s})$, indicating the “reviewer” $M_{s}$ believes that the answer $A^{j}_{i}$ is better than answer $A^{k}_{i}$ with a confidence $w^{s}$. Therefore, the answer-ranking data $\mathcal{D}$ can be defined as follows,
$\mathcal{D}=\left\{({A}_{i}^{j},{A}_{i}^{k},>,w^{s})\right\}_{i\sim\mathcal{Q},j,k,s\sim\mathcal{M}}$ (6)
@@ 2.2.1 Peer Review Stage
2.2.2 Consistency Optimization Stage
As shown in Eq 6, following the peer-review mechanism, we construct anonymous answer pairs and randomly select other LLMs as “reviewers” to evaluate both responses with a learnable confidence $w$. Next, we expect to optimize the confidence $w$ and re-rank the LLMs to be closer to human rankings. We thus propose the consistency assumption, i.e., high-level LLM can evaluate others’ answers more accurately (confidence) than low-level ones, while higher-level LLM can also achieve higher answer-ranking scores. Formally, we maximize the consistency of each LLM’s capability $w$ and score $G$ with constrained optimization as follows,
@@ 2.2.2 Consistency Optimization Stage
where $\mathbf{1}\{\cdot\}$ is the indicator function that the value is 1 when the condition is met, otherwise, it is 0. $G_{j}$ denotes the response score of model $M_{j}$, which is calculated by joint evaluation of other models. Moreover, we employ Pearson correlation [42] to measure the consistency between $w$ and $G$. Note that we only introduce this straightforward implementation to validate our idea of PiCO. Other more advanced strategies may be employed to further improve the performance.
Discussion: It is worth noting that the whole process (Eq. 6 and 7) works in an unsupervised way. The only thing we do is to adaptively assign each LLM a score that matches its abilities. An intuitive example is as follows: in a real peer-review system, if the academic level of three scholars $a$, $b$, and $c$ satisfies the following relationship, $w^{a}>w^{b}>w^{c}$. So, in the ultimate ideal scenario, the ranking of the scores submitted by these three scholars should also be, $G_{a}>G_{b}>G_{c}$. In other words, the sorting of $G$ and $w$ satisfies high consistency. On the other hand, scholars with stronger abilities (i.e., scholar $a$) evaluate $A^{b}>A^{c}$ have stronger persuasiveness, so scholar $b$ should also receive higher weighted scores $1*w^{a}$.
Reviewer Elimination Mechanism. Realizing that not all LLMs have sufficient ability to evaluate the responses of other models. We thus introduce an unsupervised elimination mechanism to remove those LLMs that have low scores. It iteratively removes the lowest-scoring LLM from the “reviewer queue” for the next consistency optimization stage, until 60% of models are eliminated. The whole process of the approach is summarized in Algorithm 1, and the details can be found in Appendix D.

### v3

# PiCO: Peer Review in LLMs based on Consistency Optimization

@@ Abstract
Existing large language models (LLMs) evaluation methods typically focus on testing the performance on some closed-environment and domain-specific benchmarks with human annotations. In this paper, we explore a novel unsupervised evaluation direction, utilizing peer-review mechanisms to measure LLMs automatically without any human feedback. In this setting, both open-source and closed-source LLMs lie in the same environment, capable of answering unlabeled questions and evaluating each other, where each LLM’s response score is jointly determined by other anonymous ones. During this process, we found that those answers that are more recognized by other “reviewers” (models) usually come from LLMs with stronger abilities, while these models can also evaluate others’ answers more accurately. We formalize it as a consistency assumption, i.e., the ability and score of the model usually have consistency. We exploit this to optimize each model’s confidence, thereby re-ranking the LLMs to be closer to human rankings. We perform experiments on multiple datasets with standard rank-based metrics, validating the effectiveness of the proposed approach. Our code is released at https://github.com/PKU-YuanGroup/PiCO.
@@ 1 Introduction
Large language models (LLMs) (Brown et al., 2020; Achiam et al., 2023; Bubeck et al., 2023; Touvron et al., 2023; Ning et al., 2025) have achieved remarkable success across a variety of real-world applications (Zhao et al., 2023; Liu et al., 2023a; Ouyang et al., 2022; Yao et al., 2023; Ning et al., 2024; Yang et al., 2024). With the increasingly widespread application of these models, there is an urgent need for an effective evaluation method to ensure that their performance and usability meet the growing demands. To assess the ability level of LLMs, a large number of evaluation benchmarks have been proposed by using some small and domain-specific datasets with human-curated labels, such as MMLU (Hendrycks et al., 2020), HELM (Liang et al., 2022), Big-Bench (Srivastava et al., 2022), GLUE (Wang et al., 2018). However, these benchmarks can only measure LLMs’ core capability on a confined set of tasks (e.g. multi-choice knowledge or retrieval questions), which fails to assess their alignment with human preference in open-ended tasks adequately (Chiang et al., 2023; Li et al., 2023a; Nakano et al., 2021). On the other hand, these evaluations may suffer from benchmark leakage issue, referring that the evaluation data is unknowingly used for model training, which can also lead to misleading evaluations (Wei et al., 2023; Zhou et al., 2023). Therefore, blindly improving scores on these public benchmarks cannot always yield a large language model that truly satisfies human requirements.
@@
For assessing human preferences, recent studies have focused on building crowdsourced battle platforms with human ratings as the primary evaluation metric. Typical platforms include Chatbot Arena (Zheng et al., 2023), MT-Bench (Zheng et al., 2023), and AlpacaEval (Li et al., 2023b). It constructs anonymous battles between chatbots in real-world scenarios, where users engage in conversations with two chatbots at the same time and rate their responses based on personal preferences. While human evaluation is the gold standard for measuring human preferences, it is exceptionally slow and costly (Zheng et al., 2023; Ning et al., 2022). In addition, adding a new LLM to the crowdsourced battle platforms also poses a cold-start issue (Chang et al., 2023). Thus, a fundamental question arises: can we construct an unsupervised LLMs evaluation system without relying on any human feedback?
@@
Actually, in real human evaluation systems, people build the human-ability hierarchy based on different empirical assumptions. For example, majority voting (Feldman, 2006; Boyer & Moore, 1991; Surowiecki, 2005) and rating voting (Allahbakhsh & Ignjatovic, 2012) methods are widely used during the decision-making process, which are based on the wisdom of the crowds (Surowiecki, 2005; Budescu & Chen, 2015; Weller, 2007) and have been proven to lead to better results than that of an individual. Moreover, in the established practice of peer-review in academic research, scholars evaluate their academic level rankings based on the consistency assumption, i.e., scholars with stronger abilities usually have stronger persuasiveness for evaluating others, and these scholars can also obtain higher achievements. This paper attempts to explore whether a similar phenomenon exists in the LLMs evaluation systems.
@@
Figure 1: The framework of PiCO. In this framework, both open-source and closed-source LLMs lie in the same environment, capable of answering unlabeled questions and evaluating each other, where each LLM’s response score is jointly determined by other anonymous ones. We assign each LLM a learnable capability weight to optimize the score ranking based on the consistency assumption, while reducing the entropy of the peer-review evaluation system. The goal is to find a final score ranking that all LLMs “agree” it.
@@
In this paper, we propose PiCO, a Peer review approach in LLMs based on Consistency Optimization. In this setting, LLMs themselves act as “reviewers”, engaging in mutual assessments to achieve comprehensive, efficient, and performance evaluations without relying on manually annotated data. This method aims to address the limitations of existing evaluation approaches and provide insights into LLMs’ real-world capabilities. As shown in Figure 1, both open-source and closed-source LLMs lie in the same environment and answer the open-ended questions from an unlabeled dataset. Then, we construct anonymous answer pairs, while randomly selecting other LLMs as “reviewers” to evaluate both responses with a learnable confidence weight $w$. Finally, we employ this weight and calculate the response scores $G$ for each LLM based on the weighted joint evaluation. It is worth noting that the whole peer-review process works in an unsupervised way, and our goal is to optimize the confidence weights $w$ that re-rank the LLMs to be closer to human rankings.
@@
We perform experiments on multiple crowdsourcing datasets with standard rank-based metrics, the results demonstrate that the proposed PiCO framework can effectively obtain a large language models’ leaderboard closer to human preferences. The contributions of this paper can be summarized as follows:
  • We explore a novel unsupervised LLM evaluation direction without human feedback, i.e., utilizing peer-review mechanisms to measure LLMs automatically. All LLMs can answer unlabeled questions and evaluate each other.
@@
  • We conduct extensive experiments on three crowdsourcing datasets with three standard rank-based metrics validating the effectiveness of the proposed PiCO approach.
Figure 2: The pipeline of the PiCO. It is mainly composed of two components: the peer-review and consistency optimization stages. Specifically, in the peer-review stage, the unlabeled dataset $\mathcal{Q}$ and the LLMs pool $\mathcal{M}$ are given. Then, we let all LLMs answer each unlabeled question to obtain the response set $\mathcal{A}$. We shuffle the set and construct anonymous answer pairs, while randomly selecting other LLMs to evaluate both responses with a learnable confidence $w$. As a result, we can obtain the answer-ranking data $\mathcal{D}$ which is a quadruple that records the partial order between two answers and the evaluator’s confidence weight. In the consistency optimization stage, we update the parameter $w$ by maximizing the consistency of each LLM’s capability and score, while re-ranking the LLMs to be closer to human rankings.
@@ 2 The Proposed Approach
2.1 Problem Definition
This paper aims to re-rank the ability of LLMs to be closer to human (ground-truth) rankings $\mathcal{R}^{*}$ in an unsupervised way (without relying on any human annotations). Specifically, we have a large language models (LLMs) pool =$\mathcal{M}=\{M_{j}\}_{j=1}^{m}$, which includes both open-source and closed-source models. Write ${M}_{1}\succ{M}_{2}$ to indicate that the LLM ${M}_{1}$ has stronger capabilities than the LLM ${M}_{2}$. Thus, we can assume that the ground-truth ranking $\mathcal{R}^{*}$ is as follows,
$\mathcal{R}^{*}:=[{M}_{1}\succ{M}_{2}\succ{M}_{3}\succ...\succ{M}_{m}].$ (1)
Assuming that the learned ranking $\hat{\mathcal{R}}$ by different evaluation methods is as follows,
@@ 2.1 Problem Definition
The goal is to learn an LLM ranking $\hat{\mathcal{R}}$ that aligns with human ranking $\mathcal{R}^{*}$ as much as possible.
@@ 2.2 Algorithm Details
The pipeline of the proposed PiCO, depicted in Figure 2, involves peer-review and consistency optimization stages. Next, we will introduce the two stages in detail.
Peer Review Stage. In our peer-review system, we consider an unsupervised LLM evaluation scenario with an unlabeled dataset $\mathcal{Q}$ consisting of $n$ open-ended questions, where $\mathcal{Q}=\{Q_{i}\}_{i=1}^{n}$. All LLMs will answer each unlabeled question to obtain the set $\mathcal{A}=\{\{A_{i}^{j}\}_{i=1}^{n}\}_{j=1}^{m}$, where $A_{i}^{j}$ is as follows,
$A_{i}^{j}=M_{j}(Q_{i})$ (3)
which infers the model $M_{j}$ response an answer $A_{i}^{j}$ with question $Q_{i}$. In addition, LLMs themselves also act as “reviewers” to evaluate other answers. Specifically, for the same question $Q_{i}\in\mathcal{Q}$, we randomly construct a battle pair $<A_{i}^{j},A_{i}^{k}>$ for review. Each battle pair will randomly assign “reviewers” to determine the winners or declare ties,
$(A_{i}^{k},A_{i}^{s},>,w^{j})=M_{j}(A_{i}^{k};A_{i}^{s}|Q_{i}).$ (4)
Under the same question $Q_{i}$, the quadruples $(A_{i}^{k},A_{i}^{s},>,w^{j})$ indicate that the “reviewer” $M_{j}$ believes that the answer $A_{i}^{k}$ is better than answer $A_{i}^{k}$ with a confidence $w^{j}$. Thus, we can collect the answer-ranking data $\mathcal{D}$ as follows,
$\mathcal{D}=\left\{({A}_{i}^{k},{A}_{i}^{s},>,w^{j})\right\}_{i\sim\mathcal{Q},j,k,M_{j}\sim\mathcal{M}},$ (5)
@@
$G_{j}=\sum_{(A_{i}^{k},A_{i}^{s},>,w^{j})\sim\mathcal{D}}\mathbf{1}\{A_{i}^{j}>A_{i}^{k}\}\cdot w^{s},$ (6)
where $\mathbf{1}\{\cdot\}$ is the indicator function that the value is 1 when the condition is met, otherwise, it is 0. We can define the LLM $M_{1}$ is better than $M_{2}$ as its score is larger, i.e., $M_{1}\succ M_{2}:=G_{1}>G_{2}$. Thus, we can re-write the learned LLM ranking $\hat{\mathcal{R}}$ as follows,
^$\hat{\mathcal{R}}:=[{G}_{3}>{G}_{1}>{G}_{2}>...>{G}_{m}].$ (7)
Thus, the goal is to learn the confidence weights $w$ to adjust the final ranking $\hat{\mathcal{R}}$ to be closer to ground-truth ranking $\mathcal{R}^{*}$.
Table 1: Validation of consistency assumption. Performance comparison of Backward, Uniform, Forward weight voting, and Consistency Optimization methods with two metrics across three datasets.
Methods MT-Bench Chatbot Arena AlpacaEval
$S(\uparrow)$ $\tau(\uparrow)$ $S(\uparrow)$ $\tau(\uparrow)$ $S(\uparrow)$ $\tau(\uparrow)$
Backward Weight 0.70 0.50 0.72 0.52 0.69 0.50
Uniform Weight 0.74 0.54 0.80 0.58 0.77 0.58
Forward Weight 0.75 0.56 0.82 0.59 0.79 0.60
Random Weight + Consistency Optimization 0.90 0.77 0.89 0.72 0.84 0.68
Validation of Consistency Assumption. First of all, we start with a toy experiment to study the role of confidence $w$ in Table 1. Specifically, we manually construct three methods: Backward Weight, Uniform Weight, and Forward Weight. That is, the ability weights of the model are respectively weighted forward ($w=[1,0.9,...,0]$), uniformly ($w=[1,1,...,1]$), and backward ($w=[0,0.1,...,1]$) according to the ground-truth human ranking. In other words, the Forward Weight means manually assigning higher weights to those models with stronger abilities, and so on for others. Then, we can calculate the response score $G_{j}$ for each model using Eq.6, and obtain the LLM ranking $\hat{\mathcal{R}}$. We measure the alignment between $\hat{\mathcal{R}}$ and $\mathcal{R}^{*}$ with Spearman’s $S(\uparrow)$ and Kendall’s $\tau(\uparrow)$ rank correlation coefficient in Table 1. Note that this is an ideal experiment, as we only use the ground-truth human ranking to validate the feasibility of our idea.
As shown in Table 1, it can be observed that the Forward Weight achieves better results than the Uniform and Backward ones in all cases, while the Backward one always achieves worse results. It validates that assigning larger weights to those models with stronger capabilities can obtain better results. In other words, those answers that are more recognized by other “reviewers” (models) usually come from LLs with stronger abilities. We formalize it as a consistency assumption, i.e., high-level LLM can evaluate others’ answers more accurately (confidence) than low-level ones, while higher-level LLM can also achieve higher answer-ranking scores, the ability and score of the model usually have consistency.
Consistency Optimization Stage. Based on this observation, we propose to maximize the consistency of each LLM’s capability $w$ and score $G$ with constrained optimization as follows,
@@
where the Pearson correlation (Sedgwick, 2012) is used to measure the consistency between $w$ and $G$. Note that we only introduce this straightforward implementation to validate our idea of PiCO. Other more advanced strategies may be employed to further improve the performance.
Discussion: It is worth noting that the whole process (Eq. 5 and 8) works in an unsupervised way. The only thing we do is to adaptively adjust the score of each LLM that match its abilities. Most importantly, we also validate the effectiveness of the proposed consistency optimization in Table 1. Specifically, we randomly initialize the ability weights and employ our consistency optimization to adjust the weight. It can be observed that the learned $w$ by our consistency optimization algorithm (Eq.8) can further improve the performance of the evaluation system, making the LLM ranking $\hat{\mathcal{R}}$ closer to human ranking $\mathcal{R}^{*}$. Another intuitive example is as follows: in a real peer-review system, if the academic level of three scholars $a$, $b$, and $c$ satisfies the following relationship, $w^{a}>w^{b}>w^{c}$. So, in the ultimate ideal scenario, the ranking of the scores submitted by these three scholars should also be, $G_{a}>G_{b}>G_{c}$. In other words, the sorting of $G$ and $w$ satisfies high consistency. On the other hand, scholars with stronger abilities (i.e., scholar $a$) evaluate $A^{b}>A^{c}$ have stronger persuasiveness, so scholar $b$ should also receive higher weighted scores $1*w^{a}$.
Reviewer Elimination Mechanism. Realizing that not all LLMs have sufficient ability to evaluate the responses of other models. We thus introduce an unsupervised elimination mechanism to remove those LLMs that have low scores. It iteratively removes the lowest-scoring LLM from the “reviewer queue” for the next consistency optimization stage, until 60% of models are eliminated. The discussion of the elimination mechanism can also be found in the Experiment 3.3.